# VizAbility: Enhancing Chart Accessibility with LLM-based Conversational Interaction

### v3

@@ 5.3.1. Comparison to a question-answering system with symbolic inputs
We compared our system to a similar system that focuses on chart reasoning. (Kim et al., 2020). Directly running the system proved challenging due to compatibility issues with outdated dependencies. A similar evaluation, using our dataset from blind participants, was conducted in prior work, revealing an overall factual accuracy rate of 16% (Kim et al., 2023b). This evaluation, however, was limited to just 245 queries from the total dataset. It excluded contextual queries and others deemed unanswerable due to ambiguous wording. Furthermore, the evaluation focused solely on questions related to bar and line charts, aligning with the system’s supported question types.
@@ 6. Evaluation: User Study with Blind People
Figure 8. System Usability Scale Survey Results, showing participant responses on a Likert scale from ’Strongly Disagree’ to ’Strongly Agree’ for various statements about the system’s frequency of use, ease of use, learning curve, integration of functions, consistency, need for technical support, and complexity, as part of the qualitative study.
@@
During the development process, we engaged with a blind participant who had prior experience using a screen reader on a daily basis. This participant provided feedback at two intermediate stages of development in what was an iterative design process. In addition to this intermediate prototype evaluation, we conducted a preliminary usability study with six additional blind/low-vision individuals. IRB approval was obtained before the commencement of the study. The stimuli used in the study was an early version of VizAbility. While the major features remain unchanged, the improvements made after the study are described in Section 6.6.
@@ 6.6. Changes After User Feedback
D3
Participant 4’s experience revealed a limitation in our initial contextual query handling: VizAbility often misinterpreted user intent due to missing chart information. For instance, it treated “temporal polarity” as a linguistic term, not a data dimension depicted in the chart. We addressed this by enriching prompts with chart tree view text (Figure 3), which led to correct interpretation of the data (e.g., ”temporal polarity” in global temperature data). This observation also led to exploring query ambiguity in general, resulting in pre-processing question refinement (Section 4.5).
@@
D4
Participant 2’s preference for automatic traversal suggested that manually pressing a series of keys can be cumbersome, especially for lengthy navigation paths. Originally, we aimed to ensure transparency and grant users control over the process. However, recognizing this issue, we decided to introduce an option for automatic traversal (see Section 4.4.4). While manual navigation might become tedious with familiarity, we kept it as a default option to accommodate varying technical proficiencies.
@@
D5
To address Participant 5’s concerns regarding the complexity of navigating the tree view, we introduced a conventional raw data table as an alternative (see 4.3). Though this addition may not significantly contribute to the novelty of our work, it underscores our dedication to creating an inclusive system.
@@
D6
The initial self-guided exploration of charts showed that most participants (P1, P2, P3, P5) struggled to ask questions beyond data-retrieval queries. Despite recognizing the value of visual, contextual, and navigation queries, participants were unaware of these query types until they were explicitly explained by the moderator. This observation, along with P2’s suggestion for help documentation and preference for interactive guidance, led to the addition of query suggestions to the system (see Section 4.5 and Section 4.6).
@@ 7.1. Limitations and Opportunities
Despite our initial aim to offer concise and informative answers, P5’s recommendation for user-adjustable response verbosity underscored the importance of user agency over designer-imposed settings. Given that speech is processed serially, the text length read by screen readers becomes a pivotal design consideration. This concern has been reiterated in prior research (Zong et al., 2022; Ault et al., 2002; w3c, 2023; Jung et al., 2022). Similarly, offering users the capability to customize node descriptions in the tree view could prove advantageous.
@@ 7.2. Need for Rigorous and Inclusive Benchmark Testing
Our evaluation was constrained not only by the four chart types but also by the limited range of questions (e.g., three query types), preventing a full assessment of VizAbility’s capabilities. For example, although we noted some visual queries necessitating reasoning over data and contextual queries seeking chart information, our limited dataset was not sufficient to capture multi-category questions requiring advanced multi-hop reasoning (Chen et al., 2020). Furthermore, our study did not evaluate situational questions related to a user’s current interest point within the tree view, which our dataset lacks. Questions dependent on understanding previous conversational context were also not explored. Considering the generative abilities of LLMs, synthetically generating these types of questions using human-created examples and advanced prompt engineering might be a viable method (Ko et al., 2023).
@@ Acknowledgements.
[currently redacted for anonymous review]

### v4

@@ 5.3.1. Comparison to a question-answering system with symbolic inputs
We compared our system to a similar system that focuses on chart reasoning (Kim et al., 2020). Directly running the system proved challenging due to compatibility issues with outdated dependencies. A similar evaluation, using our dataset from blind participants, was conducted in prior work, revealing an overall factual accuracy rate of 16% (Kim et al., 2023b). This evaluation, however, was limited to just 245 queries from the total dataset. It excluded contextual queries and others deemed unanswerable due to ambiguous wording. Furthermore, the evaluation focused solely on questions related to bar and line charts, aligning with the system’s supported question types.
@@ 6. Evaluation: User Study with Blind People
Figure 8. System Usability Scale Survey Results, showing participant responses on a Likert scale from ‘Strongly Disagree’ to ‘Strongly Agree’ for various statements about the system’s frequency of use, ease of use, learning curve, integration of functions, consistency, need for technical support, and complexity, as part of the qualitative study.
@@
During the development process, we engaged with a blind participant who had prior experience using a screen reader on a daily basis. This participant provided feedback at two intermediate stages of development in what was an iterative design process. In addition to this intermediate prototype evaluation, we conducted a preliminary usability study with six additional blind/low-vision individuals, during which we prioritized self-guided investigation over pre-formed questions to simulate real-world encounters with charts. IRB approval was obtained before the commencement of the study. The stimuli used in the study was an early version of VizAbility. While the major features remain unchanged, the improvements made after the study are described in Section 6.6.
@@ 6.6. Changes After User Feedback
D3 Participant 4’s experience revealed a limitation in our initial contextual query handling: VizAbility often misinterpreted user intent due to missing chart information. For instance, it treated “temporal polarity” as a linguistic term, not a data dimension depicted in the chart. We addressed this by enriching prompts with chart tree view text (Figure 3), which led to correct interpretation of the data (e.g., ”temporal polarity” in global temperature data). This observation also led to exploring query ambiguity in general, resulting in pre-processing question refinement (Section 4.5).
@@
D4 Participant 2’s preference for automatic traversal suggested that manually pressing a series of keys can be cumbersome, especially for lengthy navigation paths. Originally, we aimed to ensure transparency and grant users control over the process. However, recognizing this issue, we decided to introduce an option for automatic traversal (see Section 4.4.4). While manual navigation might become tedious with familiarity, we kept it as a default option to accommodate varying technical proficiencies.
@@
D5 To address Participant 5’s concerns regarding the complexity of navigating the tree view, we introduced a conventional raw data table as an alternative (see 4.3). Though this addition may not significantly contribute to the novelty of our work, it underscores our dedication to creating an inclusive system.
@@
D6 The initial self-guided exploration of charts showed that most participants (P1, P2, P3, P5) struggled to ask questions beyond data-retrieval queries. Despite recognizing the value of visual, contextual, and navigation queries, participants were unaware of these query types until they were explicitly explained by the moderator. This observation, along with P2’s suggestion for help documentation and preference for interactive guidance, led to the addition of query suggestions to the system (see Section 4.5 and Section 4.6).
@@ 7.1. Limitations and Opportunities
Despite our initial aim to offer concise and informative answers, P5’s recommendation for user-adjustable response verbosity underscored the importance of user agency over designer-imposed settings. Given that speech is processed serially, the text length read by screen readers becomes a pivotal design consideration. This concern has been reiterated in prior research (Zong et al., 2022; Ault et al., 2002; w3c, 2023; Jung et al., 2022; Jones et al., 2024). Similarly, offering users the capability to customize node descriptions in the tree view could prove advantageous.
@@ 7.2. Need for Rigorous and Inclusive Benchmark Testing
Our evaluation was constrained not only by the four chart types but also by the limited range of questions (e.g., three query types), preventing a full assessment of VizAbility’s scalability to larger datasets and more nuanced queries. For example, although we noted some visual queries necessitating reasoning over data and contextual queries seeking chart information, our limited dataset was not sufficient to capture multi-category questions requiring advanced multi-hop reasoning (Chen et al., 2020). Furthermore, our study did not evaluate situational questions related to a user’s current interest point within the tree view, which our dataset lacks. Questions dependent on understanding previous conversational context were also not explored. Considering the generative abilities of LLMs, synthetically generating these types of questions using human-created examples and advanced prompt engineering might be a viable method (Ko et al., 2023).
@@ Acknowledgements.
The authors would like to thank Jacob Ottiger for initial guidance with VizAbility’s system architecture, Sarah Zhang and Stephen Gwon for their work on synthesizing ground truths to dataset of BLV user queries, Nhung Van for helping develop the web application, and the reviewers for their insightful feedback. This research was supported in part by the Research Incentive Grant provided by Boston College.